OpenStudy (anonymous):

predict the next three numbers in this pattern 4, 2, 1, 1 1/2, 1/4

OpenStudy (anonymous):

hmm as u can see 4/2=2 2/2=1 im scared of 1 1/2 term xD i think it should be 1/2=1/2 and 1/2 / 2=1/4 @ganeshie8 ? what about 1 1/2 xD

OpenStudy (anonymous):

it may be geometric sequence with coefficient q=1/2;